A population is given by 5 8 10 15 16 20 draw all possible samples of size 2 without replacement. Show that E(x ̅ )=μ




1
Expert's answer
2021-11-22T12:56:22-0500

We have population values "5, 8, 10, 15, 16, 20," population size "N=6"

"\mu=\dfrac{5+8+10+15+16+20}{6}=37/3"


We have population values "5, 8, 10, 15, 16, 20," population size "N=6" and sample size "n=2." Thus, the number of possible samples which can be drawn without replacement is


"\dbinom{N}{n}=\dbinom{6}{2}=15""\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Sample & Sample & Sample \ mean \\
 No. & values & (\bar{X}) \\ \hline
 1 & 5,8 & 6.5 \\
 \hdashline
 2 & 5,10 & 7.5 \\
 \hdashline
 3 & 5,15 & 10 \\
 \hdashline
 4 & 5,16 & 10.5 \\
 \hdashline
 5 & 5,20 & 12.5 \\
 \hdashline
 6 & 8,10 & 9 \\
 \hdashline
 7 & 8,15 & 11.5 \\
 \hdashline
 8 & 8,16 & 12 \\
 \hdashline
 9 & 8,20 & 14 \\
 \hdashline
 10 & 10,15 & 12.5 \\
 \hdashline
 11 & 10,16 & 13 \\
 \hdashline
 12 & 10,20 & 15 \\
 \hline
 \hdashline
 13 & 15,16 & 15.5 \\
 \hdashline
 14 & 15,20 & 17.5 \\
 \hdashline
 15 & 16,20 & 18 \\
\end{array}"


The sampling distribution of the sample means.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
& \bar{X} & f & f(\bar{X}) & \bar{X}f(\bar{X})& \bar{X}^2f(\bar{X}) \\ \hline
 & 6.5 & 1 & 1/15 & 6.5/15 & 42.25/15 \\
 \hdashline
 & 7.5 & 1 & 1/15 & 7.5/15 & 56.25/15 \\
 \hdashline
 & 9 &1 & 1/15 & 9/15 & 81/15 \\
 \hdashline
 & 10 & 1 & 1/15 & 10/15 & 100/15 \\
 \hdashline
 & 10.5 & 1 & 1/15 & 10.5/15 & 110.25/15 \\
 \hdashline
 & 11.5 & 1 & 1/15 & 11.5/15 & 132.25/15 \\
 \hdashline
 & 12 & 1 & 1/15 & 12/15 & 144/15 \\
 \hdashline
 & 12.5 & 2 & 2/15 & 25/15 & 312.5/15 \\
 \hdashline
 & 13 & 1 & 1/15 & 13/15 & 169/15 \\
 \hdashline
 & 14 & 1 & 1/15 & 14/15 & 196/15 \\
 \hdashline
 & 15 & 1 & 1/15 & 15/15 & 225/15 \\
 \hdashline
 & 15.5 & 1 & 1/15 & 15.5/15 & 240.25/15 \\
 \hdashline
 & 17.5 & 1 & 1/15 & 17.5/15 & 306.25/15 \\
 \hdashline
 & 18 & 1 & 1/15 & 18/15 & 324/15 \\
 \hdashline
 Total & & 15 & 1 & 37/3 & 814/5 \\ \hline
\end{array}"




"E(\bar{X})=\sum\bar{X}f(\bar{X})=\dfrac{37}{3}"


The mean of the sampling distribution of the sample means is equal to the the mean of the population.


"E(\bar{X})=\mu_{\bar{X}}=\dfrac{37}{3}=\mu"